OpenStudy (anonymous):

GIVE POINT (-6,-2) AND A SLOPE OF 5, WRITE AN EQUATION IN SLOPE -INTERCEPT FORM

OpenStudy (anonymous):

slope intercept form is y=mx+b m=slope b=y-intercept

OpenStudy (anonymous):

sub -6 for x and -2 for y

OpenStudy (anonymous):

-2=(5)(-6)+b

OpenStudy (anonymous):

-2=-30+b 28=b y=5x+28 is your equation

OpenStudy (anonymous):

aliter u can use point slope form y-y1=m(x-x1) i.e y+2=5(x+6) or y+2=5x+30 or y=5x+30-2 y=5x+28 IN SLOPE -INTERCEPT FORM